# Emotion Classification In Software Engineering Texts: A Comparative Analysis of Pre-trained Transformers Language Models

### 19 Jan 2024 v1

@@ ABSTRACT.
conference: 46th International Conference on Software Engineering; April 2024; Lisbon, Portugal
@@ 1. INTRODUCTION
The exploration of emotion classification in software engineering texts has garnered considerable attention in recent years (novielli2023emotion, ; lin2022opinion, ), leading to the development of various emotion classification tools by researchers. Notable among these are ESEM-E by Murgia et al. (esem-e, ), EMTk by Calefato et al. (emtk, ), and SEntiMoji by Chen et al. (sentimoji, ). Chen et al.’s comparative study, utilizing emotion datasets extracted from Stack Overflow and JIRA (ortu2016emotional, ; novielli2018gold, ), affirmed the superior performance of SEntiMoji compared to EMTk and ESEM-E, solidifying its prominence in the domain. However, compared to open-domain, Pre-Trained Language Models (PTMs) rarely have been used in software engineering text for emotion classification.
@@ 2.2. Compared PTMs
We compare 6 encoder-based PTMs: BERT, RoBERTa, DeBERTa, ALBERT, CodeBERT and GraphCodeBERT. Previous research shows that BERT, RoBERTa and ALBERT work well in SE affect analysis (zhang2020sentiment, ; batra2021bert, ), while DeBERTa shows promising results in sentiment analysis (wang2022auto, ). In addition to these four models, we compare against 2 domain-specific models: CodeBERT and GraphCodeBERT – which showed promising results in various software engineering tasks (zhou2021assessing, ; karmakar2021pre, ). The models are explained below.
@@ 3. EXPERIMENTS
Error analysis: In order to identify the sources of mistakes made by the models, we perform an error analysis exclusively on the GitHub benchmark. We choose this dataset because it has undergone prior error analysis (imran2022data, ). We focus on 67 cases where all six models concurred, yet the ground truth differed (9 false positives and 58 false negatives). The distribution is as follows: Surprise 15, Joy 14, Anger 13, Sadness 11, Fear 9 and Love 5. For qualitative assessment, we employ the error categorization proposed by Novielli et al. (novielli2018benchmark, ). We apply a thematic approach. Initially, one of the authors of this paper did the error mapping using Novielli et al. (novielli2018benchmark, )’s categorization. Then, a senior undergraduate student reviewed the annotation and marked the disagreement that were resolved via discussion.
@@
In open-domain research, the incorporation of word polarity has proven beneficial for sentiment analysis (tian2020skep, ; zhang2023empirical, ; ke2020sentilare, ). We aim to investigate whether a similar enhancement in word polarity can improve emotion classification, specifically within the software engineering domain.
@@ 3.2.2. Results and Discussion
We observe that there remains a considerable amount of utterances (10/13) that are misclassified contain emojis. For example, the previously mentioned utterance, “And yes, there should be tests face-screaming-in-fear face-screaming-in-fear face-screaming-in-fear” - all models still predict incorrectly.
RQ2 Takeaway: Polarity-enhanced PTMs consistently outperform both the state-of-the-art SEntiMoji model and baseline PTMs in both datasets. The improvements are evident in micro-averaged and macro-averaged F1-scores, with substantial improvements observed for individual emotions for different models. While the findings suggest that incorporating sentiment polarity information into PTMs can enhance their performance in recognizing emotions, they do not always help to capture emotions, espectively when it is context dependant and figrurative language are involved.
@@ 4. DISCUSSION
In this section, we summarize the key lessons learned from our study and outline promising directions for future work based on our findings.
@@ 4.1. Lessons Learned
Our study provides several key lessons regarding the application of Pre-Trained Language Models for emotion classification in software engineering texts:
  • General domain PTMs like BERT, RoBERTa and DeBERTa consistently outperform specialized models like CodeBERT and GraphCodeBERT. This indicates that the pre-training objectives and textual domains hold greater relevance than task-specific tuning in emotion classification task in software engineering text.
  • The inclusion of polarity features during PTM fine-tuning consistently improves performance, highlighting the value of sentiment awareness. However, certain emotions, particularly negative ones like Anger, Fear, and Sadness, still pose challenges in terms of context understanding, figurative language, and pragmatics.
  • No single PTM excels across all metrics, necessitating careful benchmarking and tradeoff analysis.
  • Common error categories include general errors, implicit polarity, figurative language, and contextual gaps, underscore the difficulties in nuanced emotion recognition, requiring further customization to address these limitations.
  • Persistent challenges arise in handling emojis, with models struggling to interpret emotive signals. Dedicated approaches to handle emojis may mitigate this issue.
@@ 4.2. Future Directions
Building upon the lessons learned, promising avenues for future work include:
  • Research show that aspect-based sentiment analysis (absa)-enhanced PTMs outperform baseline PTMs (yang2023pyabsa, ; zhang2023empirical, ). As we have already demonostrated polarity-enhanced PTMs outperform baseline PTMs, SE-specific Absa-enhanced PTMs may further improve the classifiers.
  • Exploring multi-modal fusion of text and emojis cues in pre-training/fine-tuning step can improve interpretation of emotive expressions.
  • Employing hierarchical emotion classification and joint modeling of related polarity dimensions merits investigation. Emotion classification often treats each emotion category (e.g. joy, sadness, anger) separately. A hierarchical approach would involve first identifying broad emotional valence (positive, negative or neutral) and then classifying the specific emotion category. The hypothesis is that a coarse-to-fine hierarchical framework may improve performance by leveraging the intrinsic relationships and groupings among emotions. The study by Li et al. (li2023two, ) showed potential in this aspect.
  • In SE-specific models, pre-training with polarity-word masking, emojis, emoticons in top of general word masking merits an inspection. In general domain, this approach showed promise in sentiment analysis (zhou2020sentix, ).
  • A major contribution in this area will be establishing more benchmark datasets.
  • Exploring generative language models like GPT-4 and LLaMA for emotion detection using zero-shot and few-shot learning. Recent advancements allow synthesizing natural language responses, facilitating data augmentation and prompting techniques for improved classification (bayer2022survey, ; kocon2023chatgpt, ; koptyra2023clarin, ; nedilko2023generative, ).
@@ 5. RELATED WORK
On a large scale the related work can be divided into two parts: emotions in software engineering texts and pre-trained transformer models in software engineering tasks.
@@ 5.1. Emotions in Text-Based Software Engineering Communication
Pletea et al. explored the emotions expressed by developers during security discussions on GitHub (pletea2014security, ). Ortu et al. analyzed the emotional expressions in JIRA issue comments and their impact on collaboration and software development outcomes (ortu2016emotional, ). Novielli et al. annotated a benchmark for emotion annotation in Stack Overflow using Shaver’s (Shaver, ) emotion model (novielli2018gold, ). Calefato et al. developed a toolkit for emotion recognition from text (emtk, ). Mäntylä et al. explored the possibilities of using Valence, Arousal, and Dominance (VAD) dimensions (russell1977evidence, ) for detecting burnout and productivity (mantyla2016mining, ). Islam et al. proposed DEVA and MarValous – two machine learning methods for sensing emotions in the VAD space (islam2018deva, ; islam2019marvalous, ).
Werder et al. proposed a method for extracting emotions from GitHub repositories (werder2018meme, ). Destefanis et al. studied the measurement of affects in 370K GitHub issue comments to understand the emotional impact of commenters on the overall affect of GitHub issues (destefanis2018measuring, ). Ortu et al. conducted a mining analysis and empirical study of communication patterns in GitHub repositories (ortu2018mining, ; ortu2019empirical, ). Venigalla and Chimalakonda explored the emotions in software documentation (venigalla2021understanding, ). Neupane et al. presented an end-to-end approach for investigating emotion dynamics in software development (neupane2019emod, ).
Chen et al. utilized emojis for sentiment analysis and emotion detection in developers’ communication through deep neural networks (sentimoji, ). Rong et al. empirically studied emoji usage in software development communication, exploring frequency and diversity across different channels to understand developers’ emotional expression (rong2022empirical, ). Bleyl et al. applied a BERT-based emotion recognition model to Stack Overflow posts, aiming to identify developers’ expressed emotions (bleyl2022emotion, ). Imran et al. proposed an enhanced model based on Shaver’s approach, annotated a GitHub dataset, and introduced three data augmentation techniques to enhance Software Engineering emotion detection tool performance (imran2022data, ).
@@ 5.2. PTMs in Software Engineering Tasks
The utilization of PTMs has gained significant attention in software engineering text, offering a novel approach to address various challenges and enhance the performance of diverse tasks. One notable area of research involves the application of Lanaguage Models in code-related tasks. Researchers have explored the effectiveness of models like in understanding and generating code snippets. These studies aim to improve code completion, summarization, comprehension and documentation (ciniselli2021empirical, ; gao2023code, ; haque2022semantic, ; macneil2023experiences, ; rai2022review, ). Lanaguage models have shown promise in tasks related to software bugs and assessing code quality (ciborowska2022fast, ; ardimento2020using, ; wang2023clebpi, ; ali2023bert, ; zhou2021assessing, ; keim2020does, ). There also has been study on understanding code attentions in BERT (sharma2022exploratory, ). Comprehensive studies on sentiment analysis in software engineering text, tag recommendation in Stack Overflow and software aspect-based API review using PTMs have been studied (zhang2020sentiment, ; yang2022aspect, ; he2022ptm4tag, ).
In this study, we do comprehensive analysis of emotion analysis in software engineering text using PTMs and we also apply approaches in changing attention layer of PTMs for further improving the classification task.
@@ 6. THREATS TO VALIDY
Internal validity. Internal validity focuses on the extent to which the study’s findings can be attributed to the manipulation of the independent variable. Threats may arise from mistakes in our experiments, but we have provided a replication package and outputs for independent verifications. Another potential threat is the absence of cross-validation on smaller datasets, which we addressed by using stratified sampling for representativeness and an 80%-20% train-test split. There’s also a potential threat related to data leakage in SEntiMoji models against the GitHub benchmark dataset, but this is unlikely as SEntiMoji was pre-trained before 2019 (sentimoji, ), and most data points of the GitHub dataset came afterward. However, verification of this is beyond the scope of our study.

### 03 Feb 2024 v3

@@ Abstract.
conference: 3rd International Workshop on Natural Language-based Software Engineering; April 2024; Lisbon, Portugal
@@ 1. INTRODUCTION
The exploration of emotion classification in software engineering (SE) texts has garnered considerable attention in recent years (novielli2023emotion, ; lin2022opinion, ), leading to the development of various emotion classification tools by researchers. Notable among these are ESEM-E by Murgia et al. (esem-e, ), EMTk by Calefato et al. (emtk, ), and SEntiMoji by Chen et al. (sentimoji, ). Chen et al.’s comparative study, utilizing emotion datasets extracted from Stack Overflow and JIRA (ortu2016emotional, ; novielli2018gold, ), affirmed the superior performance of SEntiMoji compared to EMTk and ESEM-E, solidifying its prominence in the domain. However, compared to open-domain, Pre-Trained Language Models (PTMs) rarely have been used in software engineering text for emotion classification.
@@ 2.2. Compared PTMs
We compare 6 encoder-based PTMs: BERT, RoBERTa, DeBERTa, ALBERT, CodeBERT and GraphCodeBERT. Previous research shows that BERT, RoBERTa and ALBERT work well in SE affect analysis (zhang2020sentiment, ; batra2021bert, ), while DeBERTa shows promising results in sentiment analysis (wang2022auto, ). In addition to these four models, we compare against 2 domain-specific models that are widely used and proposed by Microsoft: CodeBERT and GraphCodeBERT – which showed promising results in various SE tasks (zhou2021assessing, ; karmakar2021pre, ). The models are explained below.
@@ 3. EXPERIMENTS
Error Analysis: To identify model mistakes, we conduct an error analysis. Due to space constraints, we focus exclusively on the GitHub benchmark which has gone prior error analysis (imran2022data, ). To understand where the PTMs commonly make mistake, we examine 67 cases where all six models agreed, yet the ground truth differed (9 false positives, 58 false negatives), distributed across emotions. For assessment, we employ Novielli et al.’s (novielli2018benchmark, ) error categorization, using a thematic approach. An author initially mapped errors, and a senior undergraduate student reviewed and resolved disagreements through discussion.
@@
In open-domain research, leveraging word polarity has proven effective for sentiment analysis (tian2020skep, ; zhang2023empirical, ; ke2020sentilare, ). We aim to explore its applicability to emotion classification within the SE domain.
@@ 3.2.2. Results and Discussion
We observe that there still remains a considerable amount of utterances (10/13) that are misclassified contain emojis. For example, the previously mentioned utterance, “And yes, there should be tests face-screaming-in-fear face-screaming-in-fear face-screaming-in-fear” - all models still predict incorrectly.
RQ2 Takeaway: Polarity-enhanced PTMs consistently outperform the SEntiMoji model and baseline PTMs in both datasets. Notable improvements are seen in micro- and macro-averaged F1-scores, along with substantial enhancements for individual metric across various models. While the findings indicate that integrating sentiment polarity improves PTM performance, it do not always help to capture context-dependent emotions.
@@ 4. DISCUSSION
In this section, based on our experiments, we summarize the key lessons learned and outline promising directions for future work related to emotion classification in SE.
@@ 4.1. Lessons Learned
Our study yields several key insights into the application of PTMs for emotion classification in SE texts. Firstly, general domain PTMs such as BERT, RoBERTa, and DeBERTa consistently outperform specialized models. This suggests that the pre-training objectives and textual domains carry greater relevance than task-specific tuning in the emotion classification task within SE text.
Secondly, the incorporation of polarity features during fine-tuning consistently enhances performance, emphasizing the value of sentiment awareness. However, challenges persist, especially with negative emotions like Anger and Fear, which pose difficulties in context understanding.
Thirdly, no single model excels across all emotions in all metrics, necessitating careful benchmarking and tradeoff analysis. Common error categories, such as implicit polarity, figurative language, and pragmatics, underscore the challenges in understanding contextual gaps in nuanced emotion recognition. Addressing these complexities requires further customization. For example, Imran et al. (imran2023shedding, ) addressed the issue of understanding figurative language in SE text using contrastive learning.
Lastly, persistent challenges arise in handling emojis, as models struggle to interpret emotive signals. Incorporating dedicated approaches to handle emojis may offer mitigation strategies.
@@ 4.2. Future Directions
Building on lessons learned, future work can focus on establishing more benchmark datasets, particularly emphasizing implicit and context-dependent emotions.
Investigating hierarchical emotion classification and joint modeling of related polarity dimensions merits investigation. Recognizing that emotion classification often treats each category separately, a hierarchical framework may enhance performance by first identifying broad emotional valence (positive, negative, or neutral) before classifying specific categories, as demonstrated by Li et al. (li2023two, ).
In SE-specific models, exploring pre-training with polarity-word masking, emojis, and emoticons atop general word masking is warranted, given promising results in the general domain for sentiment analysis (zhou2020sentix, ). Research also suggests that aspect-based sentiment analysis (ABSA)-enhanced PTMs outperform baseline PTMs (yang2023pyabsa, ; zhang2023empirical, ). As polarity-enhanced PTMs have already shown superiority over baseline PTMs, exploring ABSA-enhanced PTMs may further improve classifiers.
Additionally, investigating the multi-modal fusion of text and emojis cues during the pre-training/fine-tuning step could enhance the interpretation of emotive expressions. Exploring generative language models like GPT-4 and LLaMA for emotion detection using zero-shot and few-shot learning presents an intriguing direction. Recent advancements in these models enable synthesizing natural language responses, facilitating data augmentation and prompting techniques for improved classification (bayer2022survey, ; kocon2023chatgpt, ; koptyra2023clarin, ; nedilko2023generative, ). Focusing on detecting the most relevant emotions that may harm productivity, such as Frustration and Disappointment, holds merit (wrobel2013emotions, ). For example, a recent case study delved into understanding Frustration in OSS text using generative models in zero-shot learning (imran2023uncovering, ).
@@ 5. RELATED WORK
On a large scale the related work can be divided into two parts: emotions in SE texts and PTMs in SE tasks.
@@ 5.1. Emotions in Text-Based Software Engineering Communication
Pletea et al. explored the emotions expressed by developers during security discussions on GitHub (pletea2014security, ). Ortu et al. analyzed the emotional expressions in JIRA issue comments and their impact on collaboration and software development outcomes (ortu2016emotional, ). Novielli et al. annotated a benchmark for emotion annotation in Stack Overflow using Shaver’s (Shaver, ) emotion model (novielli2018gold, ). Calefato et al. developed a toolkit for emotion recognition from text (emtk, ). Mäntylä et al. explored the possibilities of using Valence, Arousal, and Dominance (VAD) dimensions (russell1977evidence, ) for detecting burnout and productivity (mantyla2016mining, ). Islam et al. proposed MarValous – a machine learning method for sensing emotions in the VAD space (islam2019marvalous, ). Werder et al. proposed a method for extracting emotions from GitHub repositories (werder2018meme, ). Destefanis et al. studied the measurement of affects in 370K GitHub issue comments to understand the emotional impact of commenters on the overall affect of GitHub issues (destefanis2018measuring, ). Ortu et al. conducted a mining analysis on finding emotional state of contributors in GitHub repositories (ortu2018mining, ). Venigalla et al. explored the emotions in software documentation (venigalla2021understanding, ). Neupane et al. investigated emotion dynamics in software development (neupane2019emod, ).
Chen et al. utilized emojis for sentiment analysis and emotion detection in developers’ communication through deep neural networks (sentimoji, ). Rong et al. empirically studied emoji usage in software development communication, exploring frequency and diversity across different channels to understand developers’ emotional expression (rong2022empirical, ). Bleyl et al. applied a BERT-based emotion recognition model to Stack Overflow posts, aiming to identify developers’ expressed emotions (bleyl2022emotion, ). Imran et al. proposed an enhanced model based on Shaver’s approach for software engineering text and introduced three data augmentation techniques to enhance Software Engineering emotion detection tool performance (imran2022data, ). Tulli et al. (tulili2023burnout, ) did a literature review on burnout in software engineering.
@@ 5.2. PTMs in Software Engineering Tasks
The utilization of PTMs has gained significant attention in software engineering text, offering a novel approach to address various challenges and enhance the performance of diverse tasks. One notable area of research involves the application of PTMs in code-related tasks. Researchers have explored the effectiveness of models like in understanding and generating code snippets. These studies aim to improve code completion, summarization, comprehension and documentation (ciniselli2021empirical, ; gao2023code, ; haque2022semantic, ; macneil2023experiences, ; rai2022review, ). Lanaguage models have shown promise in tasks related to software bugs and assessing code quality (ciborowska2022fast, ; ardimento2020using, ; wang2023clebpi, ; ali2023bert, ; zhou2021assessing, ; keim2020does, ). There also has been study on understanding code attentions in BERT (sharma2022exploratory, ). Researchers also conduct studies on benchmarking PTMs on tag recommendation in Stack Overflow text, sentiment analysis in SE text and software aspect-based API review text (zhang2020sentiment, ; yang2022aspect, ; he2022ptm4tag, ).
In this study, we do comprehensive benchmark analysis of emotion classification task in SE text using PTMs and we also apply approaches in changing attention layer for further improving the performances.
@@ 6. THREATS TO VALIDY
Internal validity. Internal validity focuses on the extent to which the study’s findings can be attributed to the manipulation of the independent variable. Threats may arise from mistakes in our experiments, but we have provided a replication package and outputs for independent verifications. Another potential threat is the absence of cross-validation on smaller datasets, which we addressed by using stratified sampling for representativeness and an 80%-20% train-test split. There’s also a potential threat related to data leakage in SEntiMoji models against the GitHub benchmark dataset, but this is unlikely as SEntiMoji was pre-trained before 2019 (sentimoji, ), and data points of the GitHub dataset came afterward (imran2022data, ). However, verification of this is beyond the scope of our study given that not all pre-trained data is released by the authors (sentimoji, ).